40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4299     Entry time: Tue Feb 15 11:53:02 2011
Author: kiwamu 
Type: Summary 
Category: General 
Subject: today's missions 

I think this week is going to be an "alignment week".

The goal is to get a good alignment on X arm for both the green and the IR beam in order to reduce a2l couplings.

 

Today's missions are :

 - fixing the oplev channel names (see here)

 - fixing the oplev gain issue (see here)

 - engage the oplev servos

 - f2p adjustment

 - make a realtime lockin model for the f2p measurement and the dithering technique

 - alignment of the MC incident beam  (because we installed a new PMC this morning)

 - manual alignment of the IR beam by steering PZT1 and PZT2 (this procedure will be replaced by an automatic way soon)

 - bounce roll filters (see here)

 

ELOG V3.1.3-